# Harder or Different? Understanding Generalization of Audio Deepfake Detection

### v1

@@ 1 Introduction
Recent years have seen tremendous advances in machine learning (ML). One area in which progress has been particularly impressive is that of text-to-speech (TTS) synthesis. It is now possible to generate high-quality, convincing speech signals which mimic closely the voice identity [1, 2, 3]. Numerous online services, e.g. [4, 5, 6], enable this technology to be used by anyone, even those without any technical expertise. While it has plentiful legitimate applications, its accessibility has also given rise to serious threats, e.g. fraud, misinformation, or defamation [7, 8, 9, 10].
@@
A potential solution lies in the ML-driven detection of such deepfakes using, for example, binary classifiers to discriminate between genuine/bonafide and AI-generated speech. The field has witnessed a surge in research, from the development of new detection models [11, 12, 13, 14, 15, 16] to the creation of extensive datasets [17, 18, 19, 20, 21, 22, 23]. Most notably, initiatives such as ASVspoof [24, 25, 26] which were launched to benchmark competing detection solutions, seemingly show impressive progress; lower and lower state-of-the-art error rates are reported on a regular basis [15, 16, 27]. However, reliability in real-world scenarios often remains untested, while there are reports that generalisation to out-of-domain scenarios is wanting. The limited ability to generalize to deepfakes generated using new attack algorithms, or even algorithms that are simply different to those used to create training data, has been and continues to be a source of major concern [17].
@@
Despite generalisation having been a focal point of related research for almost a decade, we remain far from practical deepfake detection solutions which generalise to attacks and acoustic conditions seen in the wild, c.f. Figure 1. We assume that one key to unlocking progress is an understanding of why current detection solutions fail to generalize. The answer can be attributed to two main factors. First, for want of a better term, ‘hardness’: the increasing sophistication of speech deepfakes may make them inherently more challenging to detect. Second, ‘difference’: the characteristics used by a detection model to discriminate between bonafide speech and deepfakes may not generalize across different attack algorithms. This implies that detection failures might not result from a lack of detection model capacity, but from fundamentally different deepfake characteristics and features. Thus, the issue of generalization may be due to either ‘hardness’ or ‘difference’, or a combination thereof.
@@
Contributions. We report a means to decompose the performance discrepancy between training and out-of-domain test data into two components: ‘hardness’ and ‘difference’. We report a study, performed using four different detection models and the ASVspoof 2019 and ASVspoof 2021 datasets, which shows that:
@@
  • When randomly degrading input in the ASVspoof 2019 database, performance is significantly diminished, which we show can be attributed mainly to ‘hardness’.
@@
  • In contrast, for the ASVspoof 2021 deepfake (DF) dataset which contains data collected from other sources, degraded detection performance can be attributed almost exclusively to ‘difference’. This also holds for the In-the-Wild database [17].
These findings suggest that efforts to extend model capacity, while beneficial in the case of in-domain benchmarking, are insufficient and might even be detrimental to the pursuit of generalisable detection solutions. We argue that this calls for the re-focusing of research effort to better understand and address the ‘difference’ between deepfake techniques in the wild.
@@ 2 Related Work
A significant research in speech deepfake detection was performed in the scope of ASVspoof challenges [24, 28, 25, 26]. This body of work has established a benchmark that subsequent work has taken as a reference point to assess performance [11, 12, 13, 14, 15, 16]. However, recent observations highlight that promising results obtained using ASVspoof databases do not necessarily translate to reliable detection in real-world scenarios [29, 17]. This finding has prompted the creation of new datasets designed to address the performance gap [23]. Research has also explored the detection of so-called partial-spoofs, where bonafide utterances are segmented and then concatenated with content generated using text-to-speech synthesis [20]. Other directions include the investigation of deepfake detection for singing voices [22]. Despite progress, generalisation [30] issues and robustness [31] remain in focus, new attacks continue to emerge and so the challenge is still as great as ever. Whereas specific studies of generalisation have been reported in other fields, e.g. computer vision and object detection [32, 33], a similar comprehensive analysis of the underlying challenges in speech deepfake detection is lacking. The related work in computer vision serves as inspiration and as a methodological foundation for the study reported in this paper.
@@ 3 Methodology
Let $A$ represent a learning rule, encompassing model architecture, hyperparameters, and training configuration. After training on the training partition of database $D$, we evaluate the performance $p$ on the test partition of database $D^{\prime}$. Training and test partitions are always disjoint, but may or may not be sourced from the same database, representing either an in-domain (i.e., $D\to D)$ or out-of-domain (i.e., $D\to D^{\prime})$ evaluation scenario, respectively. Performance is assessed using the Equal Error Rate (EER) which, despite well-known shortcomings, is often used when comparing independent spoofing detection solutions [28, 26].
To gain insights into performance deteriorations stemming from generalisation issues, we compare in-domain performance $p^{A}_{D\rightarrow D}$ to out-of-domain performance $p^{A}_{D\rightarrow D^{\prime}}$ for different detection systems and datasets. To illustrate the idea, we aim to understand why models trained on the data from one database exhibit better detection performance when tested on data from the same database (in-domain) compared to data from a different database (out-of-domain). This performance is expressed in terms of the performance gap
$p^{A}_{D\rightarrow D}-p^{A}_{D\rightarrow D^{\prime}}.$ (1)
which is the difference in detection performance for in-domain and out-of-domain scenarios. It can be decomposed as follows.
$\underbrace{p^{A}_{D\rightarrow D}-p^{A}_{D\rightarrow D^{\prime}}}_{\text{Performance gap}}$ (2)
$=$ $\underbrace{p^{A}_{D\rightarrow D}-p^{A}_{D^{\prime}\rightarrow D^{\prime}}}_{\text{Hardness gap}}+\underbrace{p^{A}_{D^{\prime}\rightarrow D^{\prime}}-p^{A}_{D\rightarrow D^{\prime}}}_{\text{Difference gap}}$
Thus, we break down the performance gap into two components: hardness gap and difference gap. If the performance gap stems primarily from inherently more challenging test data, such as more sophisticated deepfakes, i.e. attacks that produce high-quality speech signals with minimal artefacts, then we anticipate a more significant ‘hardness gap’. On the other hand, if deepfakes contained in the test data exhibit a similar level of artefacts but differ in acoustic conditions compared to the training data, then the performance gap should be attributed to a greater ‘difference’ gap. We expect the same in the case that deepfakes contained in the test data are no more ‘hard’ to detect, if only learning was performed using similar deepfakes among the training data, hence the distinction between what is ‘hard’ and what is ‘different’. If the ”performance gap” is primarily due to ”difference”, then enhancing model capabilities may not address the underlying issue.
@@ 4 Evaluation
We use four learning models from related work in our experiments: The LCNN model [11], the RawNet2 [34], Whisper-DF [16] and SSL-W2V2 [35]. These models represent the most common approaches to speech deepfake detection, and were state-of-the-art at the time of their publication. When training these models, we use a batch sizes of 16, the Adam optimizer with a learning rate of $10^{-3}$, and allow up to $500$ epochs. To prevent overfitting, we use aggressive early stopping with a delta of $0.5\%$ train EER and a patience of $1$ epoch. Additionally, we remove leading and trailing silence from all data, as models tend to overly rely on non-informative features such as silence [29].
@@ 4.1 Databases
We utilize the following databases:
ASVspoof 2019 LA [28]: We use ASVspoof 2019 LA for both training and in-domain testing. It has three disjoint partitions: training, development and evaluation. Importantly, the evaluation partition includes out-of-domain data, namely, attacks not seen during the training and development phases. Since our focus is on using ASVspoof 2019 LA for training and in-domain testing, we disregard these partitions and reorganize the entire dataset into an $80\%$ training and $20\%$ in-domain testing split.
ASVspoof 2021 [36]: To evaluate with out-of-domain data, we utilize samples from the ASVspoof 2021 challenge. This includes the ‘Logical Access’ (LA) dataset and the ‘Deepfake’ (DF) dataset. Compared to ASVspoof 2019, these datasets pose additional challenges. The ASVspoof 2021 LA dataset simulates more realistic channel conditions, with all audio files being processed through real telephony systems, including voice-over-internet-protocol (VoIP) systems and public switched telephone networks (PSTN) [26]. Consequently, detection models must handle compression artifacts, packet losses, and other related issues. On the other hand, the ASVspoof 2021 DF dataset introduces new audio data, specifically, previously unseen deepfakes from the Voice Conversion Challenges of 2018 and 2020 [37, 38]. Additionally, lossy codecs are used for media storage to closely mimic realistic scenarios.
In-the-Wild [17]: This dataset includes both synthesized and genuine audio for 58 politicians and public figures, obtained from social networks and video platforms. It encompasses 20.8 hours of authentic and 17.2 hours of fake audio. This dataset poses a significant challenge, and models trained on ASVspoof 2019 have shown inadequate generalization on it.
@@ 4.2 Experiments
Audio Degradation. In our initial experiment, we deliberately lower the quality of the test data by aggressively truncating input length. In this scenario, with respect to Equation 2, $D$ denotes the original ASVspoof 2019 dataset with $8$ seconds (s) of input, while $D^{\prime}$ represents the same data where each sample is randomly truncated to $0.25s$. For models that require a minimum audio length (such as WhisperDF and SSL-W2V2), the truncated input is looped. Detailed outcomes are provided in Section 5.1.
Unseen databases. The second, third, and fourth experiments evaluate the performance gap between ASVspoof 2019 LA and new databases: ASVspoof 2021 LA, DF, and In-the-Wild, respectively. These scenarios represent progressively increasing complexity levels. While ASVspoof 2021 LA includes the same attacks [^1] In this context, an ‘attack’ refers to a distinct Text-to-Speech (TTS) synthesis algorithm. Introducing a new and unseen attack, therefore, means presenting a new and previously unencountered category of audio spoof or audio deepfake to the model. as ASVspoof 2019 but incorporates lossy compression and telephone encodings, ASVspoof 2021 DF introduces partially new attacks (sourced from the voice conversion challenge databases [37, 38]). Lastly, the In-the-Wild dataset consists entirely of new, unseen, and potentially high-quality attacks due to its recency compared to ASVspoof 2019. The results of these experiments are presented in Section 5.2 and Section 5.3.
@@ 5 Results
5.1 Audio Degradation
Table 1: Decomposition of the performance gap into ‘hardness’ and ‘difference’ for the ASVspoof 2019 dataset, when input length is truncated for test data. A significant portion of the performance gap can be attributed to ‘hardness’ (blue). Surprisingly however, for some models, there also is a rather significant contribution from ‘difference’ (red). Results presented as mean EER ± standard deviation across five individual trials.
Model Input (s) Perform. Hardn. Diff.
Train Test Gap Gap Gap
LCNN 8.0 0.25 $17.7\pm 0$ $13.4\pm 2$ $4.2\pm 2$
RawNet2 8.0 0.25 $33.4\pm 4$ $30.4\pm 3$ $3.0\pm 5$
SSL W2V2 8.0 0.25 $23.1\pm 2$ $11.8\pm 2$ $11.2\pm 3$
WhisperDF 8.0 0.25 $25.1\pm 1$ $8.4\pm 1$ $16.6\pm 1$
Recall that as per Equation 2, the ‘performance gap’ is the sum of the ‘hardness gap’ and ‘difference gap’. Table 1 illustrates this breakdown of performance degradation due to reduced input length. We observe that the main reason for this is the ‘hardness gap’, as marked in blue. As expected, diminishing the amount of information available to the model inherently increases the problem’s ‘hardness’. The RawNet2 model experiences the most pronounced effect: a $33.4\%$ increase in EER, of which $30.4\%$ EER can be attributed to increased ‘hardness’, while only $3.0\%$ EER accord to the ‘difference’ in conditions. Similarly, the LCNN model’s performance deteriorates by $17\%$ EER, of which $13\%$ EER pertain to an increase in ‘hardness’. Larger, pre-trained models like SSL W2V2 and WhisperDF, exhibit a less pronounced sensitivity to the ‘hardness gap’, but are still significantly affected. Interestingly, there also is a significant contribution to ‘difference’. This is discussed in section 6.
5.2 ASVspoof 2021 LA
Table 2 details the results for the ASVspoof 2021 LA dataset, which contains the same audio files as ASVspoof 2019, but which have been altered significantly through the use of encoding and compression techniques. Surprisingly, contrary to expectations these modifications would increase the dataset’s difficulty, this findings reveal a different scenario. Although performance deterioration is less pronounced than in the cases of input data reduction, the primary factor for the performance gap is attributed to ‘difference’. Notably, the LCNN model, the smallest among the four analyzed, shows a blend of both hardness and difference in its performance decline. In contrast, larger models predominantly link their diminished performance to ‘difference’. We will further discuss the implication of this observation on model generalization in Section 6.
Table 2: Decomposition of the performance gap when training on ASVspoof 2019 and evaluating on ASVspoof 2021 LA. For most models, the performance gap can be attributed solely to ‘difference’ (red).

### v3

@@ 1 Introduction
Recent years have seen tremendous advances in machine learning (ML). One area in which progress has been particularly impressive is that of text-to-speech (TTS) synthesis. It is now possible to generate high-quality, convincing speech signals which mimic closely the voice identity of specific individuals [1, 2]. Numerous online services, e.g. [3, 4, 5], enable this technology to be used by anyone, even those without any relevant technical expertise. While it has plentiful legitimate applications, its accessibility has also led to serious threats, e.g. fraud, misinformation, and defamation [6, 7, 8, 9].
@@
A potential solution lies in the ML-driven detection of such deepfakes using, for example, binary classifiers to discriminate between genuine/bonafide and AI-generated speech. The field has witnessed a surge in research, from the creation of extensive datasets [10, 11, 12, 13, 14, 15, 16] to the development of new detection models [17, 18, 19, 20, 21, 22]. Most notably, initiatives such as ASVspoof [23, 24, 25] which were launched to benchmark competing detection solutions, seemingly show impressive progress; lower and lower state-of-the-art error rates are reported on a regular basis [21, 22, 26]. However, reliability in real-world scenarios often remains untested, while there are reports that generalisation to out-of-domain scenarios is wanting. The limited ability to generalize to deepfakes generated using new attack algorithms, or even algorithms that are simply different to those used to create training data, has been and continues to be a source of major concern [10].
@@
Despite generalisation having been a focal point of related research for almost a decade, we remain far from practical deepfake detection solutions which generalise to attacks and acoustic conditions seen in the wild, c.f. Figure 1. Key to unlocking progress is an understanding of why current detection solutions fail to generalize. The answer can be attributed to two main factors. First, for want of a better term, ‘hardness’: the increasing sophistication of speech deepfakes may make them inherently more challenging to detect. Second, ‘difference’: the characteristics used by a detection model to discriminate between bonafide speech and deepfakes may not generalize across different attack algorithms. This implies that detection failures might not result from a lack of detection model capacity, but from fundamentally different deepfake characteristics. Thus, the issue of generalization may be due to either ‘hardness’ or ‘difference’, or a combination thereof.
@@
Contributions. We report a means to decompose the performance discrepancy between training and out-of-domain test data into two components: ‘hardness’ and ‘difference’. We report a study, performed using four different detection models and the ASVspoof 2019 and ASVspoof 2021 datasets, which shows the following.
@@
  • When using truncated utterance sub-segments (rather than the full utterance) selected from the ASVspoof 2019 database, performance is substantially diminished, which we show can be attributed mainly to ‘hardness’.
@@
  • In contrast, for the ASVspoof 2021 deepfake (DF) dataset which contains data collected from multiple sources, degraded detection performance can be attributed almost exclusively to ‘difference’. This observation also holds for the In-the-Wild database [10].
These findings suggest that efforts to extend model capacity, while beneficial in the case of in-domain benchmarking, are insufficient and might even be detrimental to the pursuit of generalisable detection solutions. We argue that this calls for the re-focusing of research effort to better understand and address the ‘difference’ between deepfakes seen in the wild.
@@ 2 Related Work
A substantial volume of research in speech deepfake detection was performed in the scope of ASVspoof challenges [23, 27, 24, 25]. This body of work established a benchmark which was used in subsequent work as a reference point to assess performance [17, 18, 19, 20, 21, 22]. However, recent observations highlight that promising results obtained using ASVspoof databases do not necessarily translate to reliable detection in real-world scenarios [28, 10]. This finding has prompted the creation of new datasets designed to address the performance gap [16]. Research has also explored the detection of so-called partial-spoofs, where bonafide utterances are segmented and then concatenated with content generated using text-to-speech synthesis [13]. Other directions include the investigation of deepfake detection for singing voices [15]. Despite progress, generalisation [29] and robustness [30] remain in focus. New attacks continue to emerge, hence the challenge is as great as ever. Whereas specific studies of generalisation have been reported in other fields, e.g. computer vision and object detection [31, 32], a similar comprehensive analysis of the underlying challenges in speech deepfake detection is lacking. The related work in computer vision serves as inspiration and as a methodological foundation for the study reported in this paper.
@@ 3 Methodology
Let $D$ and $D^{\prime}$ be speech databases (Section 4.1), each of which is partitioned into a training and test set. Model performance is evaluated using the Equal Error Rate (EER) [25, 27]. Define in-domain model performance $D_{\rightarrow}D$ as the EER for the test partition of $D$ after training on the training partition of $D$. Similarly, out-of-domain model performance $D_{\rightarrow}D^{\prime}$ denotes the EER when the model is trained using the training partition of $D$ and evaluated using the test partition of $D^{\prime}$.
To gain insights into the generalization performance of a given model, we analyze the performance gap:
$D_{\rightarrow}D-D_{\rightarrow}D^{\prime},$ (1)
which represents the gap in detection performance for in-domain $D_{\rightarrow}D$ and out-of-domain $D_{\rightarrow}D^{\prime}$ scenarios. It can be decomposed into a ‘hardness’ and ’difference’ components:
$\underbrace{D_{\rightarrow}D-D_{\rightarrow}D^{\prime}}_{\text{Performance gap}}$ (2)
$=$ $\underbrace{D_{\rightarrow}D-D^{\prime}_{\rightarrow}D^{\prime}}_{\text{Hardness gap}}+\underbrace{D^{\prime}_{\rightarrow}D^{\prime}-D_{\rightarrow}D^{\prime}}_{\text{Difference gap}}$
These components can be interpreted as follows:
  • The hardness gap is a measure of the relative difficulty in terms of the in-domain detection performance $D_{\rightarrow}D$ and $D^{\prime}_{\rightarrow}D^{\prime}$.
  • The difference gap reflects the difference in detection performance when a model trained using database $D$ is then tested using database $D^{\prime}$.
When analyzing the performance gap using Equation 2, two outcomes are possible. First, a larger ‘hardness gap’ indicates that $D^{\prime}$ is inherently more difficult than $D$. This difficulty may arise from the presence of higher-quality deepfakes in $D^{\prime}$ which contain fewer artefacts and which are hence harder to detect.
A second possible outcome is a larger ‘difference gap’, which suggests that the anti-spoofing model does not generalize well from $D$ to $D^{\prime}$, even though $D^{\prime}$ is not inherently more challenging. Put differently, if the ‘difference gap’ is large, then the ‘hardness gap’ must be small, which means that $D_{\rightarrow}D$ and $D^{\prime}\rightarrow D^{\prime}$ are equally challenging. Thus, poor results in $D\rightarrow D^{\prime}$ are not related to the complexity of $D^{\prime}$, and more the result of the model being overfitted to $D$. Learned artefacts may even stem from shortcuts, as described in [28], or other features which, while highly informative for $D$, do not transfer to $D^{\prime}$.
By analyzing and decomposing the performance gap as described above, we can gain insights into whether poor performance in $D_{\rightarrow}D^{\prime}$ is simply the result of $D^{\prime}$ being more challenging, or whether the two data distributions $D$ and $D^{\prime}$ are simply too different for the model to generalize.
@@ 4 Evaluation
We report an evaluation performed using four different detection models: an LCNN model [17]; RawNet2 [33]; Whisper-DF [22]; SSL-W2V2 [34]. These models represent the most common approaches to speech deepfake detection, and were state-of-the-art at the time of their publication. All models were trained using a batch size of 16, the Adam optimizer with a learning rate of $10^{-3}$, and for up to $500$ epochs. To prevent overfitting, we use aggressive early stopping with a hyper-parameters $\delta=5\cdot 10^{-3}$ train EER and a patience of $1$ epoch. Additionally, we remove leading and trailing silence from all data, to avoid the use of non-speech shortcuts reported in [28].
@@ 4.1 Databases
Experiments were performed using the following three databases, each of which is partitioned into training (80%) and test (20%) sets.
ASVspoof 2019 LA [27]: We use ASVspoof 2019 LA for both training and in-domain testing. It has three disjoint partitions: training, development and evaluation. Importantly, the evaluation partition includes out-of-domain data, namely, attacks not seen during training. Since our focus is on using ASVspoof 2019 LA for training and in-domain testing, we disregard these partitions and reorganize the entire dataset into an $80\%$ training and $20\%$ in-domain testing split.
ASVspoof 2021 [35]: Out-of-domain evaluation was performed using ASVspoof 2021 databases, specifically the ‘Logical Access’ (LA) dataset and the ‘Deepfake’ (DF) dataset. ASVspoof 2021 LA data contains simulated channel variability, with all audio files being processed through real telephony systems, including voice-over-internet-protocol (VoIP) systems and public switched telephone networks (PSTN) [25]. Consequently, detection models must handle compression artifacts, packet loss, and other related issues. On the other hand, the ASVspoof 2021 DF dataset contains deepfakes sourced from the 2018 and 2020 Voice Conversion Challenge databases [36, 37]. A subset of the DF dataset also contains additional variability stemming from the application of various lossy media storage codecs.
In-the-Wild [10]: This dataset includes both synthesized and genuine audio for 58 politicians and public figures, obtained from social networks and video platforms. It encompasses 20.8 hours of authentic and 17.2 hours of fake audio. Models trained using ASVspoof 2019 have been shown to generalise poorly to in-the-wild data [10].
@@ 4.2 Experiments
Audio Degradation – In our initial experiment, we deliberately lower the quality of the test data by aggressively truncating utterance duration. In this scenario, with respect to Equation 2, $D$ denotes the original ASVspoof 2019 dataset with $8$ seconds (s) of input, while $D^{\prime}$ represents the same data where each sample is randomly truncated to $0.25s$. For models that require a minimum audio length (such as WhisperDF and SSL-W2V2), the truncated input is appropriately concatenated after truncation in order to produce utterances of the required duration, c.f. Section 5.1.
Unseen databases – The second, third, and fourth experiments evaluate the performance gap between ASVspoof 2019 LA and other databases: ASVspoof 2021 LA, DF, and In-the-Wild. The ASVspoof 2021 LA database includes the same attacks [^1] In this context, an ‘attack’ refers to a distinct Text-to-Speech (TTS) synthesis algorithm. Introducing a new and unseen attack thus involves the inclusion of a new and previously unencountered TTS synthesis algorithm. as the ASVspoof 2019 database, but incorporates lossy compression and telephone encodings. The ASVspoof 2021 DF database includes a broader variety of attacks sourced from the voice conversion challenge databases [36, 37]. Lastly, the In-the-Wild dataset consists of unknown attacks which likely include some which are different to those in the ASVspoof databases. The results of these experiments are presented in Section 5.2 and Section 5.3.
@@ 5 Results
5.1 Experiment 1: Audio Degradation
Table 1: Decomposition of the performance gap into ‘hardness’ and ‘difference’ for $D$ = ASVspoof 2019, and $D^{\prime}$ = ASVspoof 2019 with input length truncated to 0.25s. A substantial portion of the performance gap can be attributed to ‘hardness’ (blue). Surprisingly, for some models, there also is a rather substantial contribution from ‘difference’ (red). Results are presented as mean EER ± standard deviation across five individual trials.
Model Performance Hardness Difference
Gap Gap Gap
LCNN $17.7\pm 0$ $13.4\pm 2$ $4.2\pm 2$
RawNet2 $33.4\pm 4$ $30.4\pm 3$ $3.0\pm 5$
SSL W2V2 $23.1\pm 2$ $11.8\pm 2$ $11.2\pm 3$
WhisperDF $25.1\pm 1$ $8.4\pm 1$ $16.6\pm 1$
Recall that as per Equation 2, the ‘performance gap’ is the sum of the ‘hardness gap’ and ‘difference gap’. Table 1 illustrates this breakdown of performance degradation due to reduced input length. We observe that the main reason for this is the ‘hardness gap’, as marked in blue. As expected, diminishing the amount of information available to the model inherently increases the problem ‘hardness’. The most pronounced effect is seen for the RawNet2 model: a $33.4\%$ increase in EER, of which $30.4\%$ EER can be attributed to increased ‘hardness’, while only $3.0\%$ EER can be attributed to the ‘difference’ in conditions. Similarly, performance for the LCNN model deteriorates by $17\%$ EER, of which $13\%$ EER pertains to an increase in ‘hardness’. Larger, pre-trained models like SSL W2V2 and WhisperDF, exhibit less sensitivity to the ‘hardness gap’, but are still substantially affected. Interestingly, there is also a substantial contribution from the ‘difference’.
5.2 Experiment 2: ASVspoof 2021 LA
Table 2 details results for the ASVspoof 2021 LA dataset. Even if the differences in compression and encoding are expected to compound the challenge, our experiments indicate otherwise: although performance deterioration is less pronounced than for input data reduction, the primary factor for the performance gap is attributed to ‘difference’. Notably, for the LCNN model, the smallest among the four analyzed, the performance gap can also be attributed to increased hardness. For the other, larger models, it seems that the presence of compression and encoding variation does not increase hardness.
Table 2: Decomposition of the performance gap where $D$ = ASVspoof 2019 and $D^{\prime}$ = ASVspoof 2021 LA. For most models, the performance gap can be attributed solely to ‘difference’ (red).